Nthanda’s budget line relating to good X and good Y has intercepts of 50 units of good X and 20 units of good Y. Suppose the price of good X is K12
a) What is Nthanda’s income (1.5 Marks)
b) What is the price of good Y? (1.5 Marks)
c) Suppose the price of good X was to fall to K10, what effect would this have on the budget line?
1
Expert's answer
2020-10-02T06:58:14-0400

Nthandas income

="P_x\times Q_x+ P_y\times Q_y"


When "Y=0,X =50 units"


="K12\times50+P_y\times0"

"K600" is total income


"B" price of good Y?

"=\frac{total income}{maximum quantity of Y that can be purchased}=\frac{K600}{20}" ="K600"


"C." Effect of the budget line

Decrease in price of good X will have anticlockwise movement of budget line .After decrease in price of X to K10, K600/K10= 60 units of X can be purchased